Grade 7Algebra

An angle is half of its supplement.Find the measure of the angle.

Shilpa
6 Years agoGrade 7

1 Answer

Vikas TU
6 Years ago
Let x = the angle
180 - x = its supplement since supplementary angles add up to 180 degrees.
180 - x = 1/2 x
2(180 - x) = 2(1/2) x
360 - 2x = x
360 = 3x
120 = x
The angle is 120 degrees, and its supplement is 60 degrees.